OpenStudy (anonymous):

Help Pleasee !! Medal && Fan (: Would You Rather Go On A Beach Vacation, A City Vacation, or Stay Home ?? Beach - 212 City - 120 Home - 113 A. Is The Survey Question Biased? Why Or Why Not ? B. Based On The Survey, What Are The Odds That A Person Will Want To Stay Home For The Vacation? C. Based On The Survey, What Is The Probability That A Person Will Travel For Vacation?

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

i dont get the q

OpenStudy (rajee_sam):

A. The survey question is not biased because it does not offer the person any incentive to choose either one of them. They are just asking their choice of vacation. B. odds of an event = ratio of the P (Event happening) to P ( Event not happening) odds of staying home is 113 : 332 = 113/332 C. P (Travel for Vacation) = 332/445 332 - vacation involving travel. 445 - total people surveyed.